### v1

# Postmerger: a new and dominant contribution to the gravitational-wave background from binary neutron stars

Abstract
The stochastic gravitational-wave background (SGWB) generated by the inspiral and merger of binary neutron stars is traditionally modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for the most massive binaries, it is not what is expected in general, where a remnant is produced and may survive for up to hundreds of milliseconds and radiate an amount of energy that is significantly larger than that lost during the whole inspiral. To account for this additional contribution to the SGWB, we consider a waveform model that includes both the inspiral and the postmerger emission. We show for the first time that for a large set of parameterized equations of state compatible with observational constraints, there is typically five times more power spectral density in the SGWB from the postmerger emission than from the inspiral one, leading to a normalized GW energy density $\Omega_{\rm GW}\simeq 10^{-10}-10^{-9}$. This power is predominantly located in the $1-2\,{\rm kHz}$ frequency range, hence distinct from that associated with the inspiral only. We discuss the significantly enhanced detectability of the SGWB with special attention to third-generation detectors, such as the Einstein Telescope and Cosmic Explorer, and show how it depends on the signal-to-noise ratio of foreground binaries and on the metastable remnant survival time. Interestingly, even a non-detection of the high-frequency part of the SGWB could provide valuable constraints on the remnant lifetime, offering novel insights into the postmerger dynamics and into the equation of state of nuclear matter.
I Introduction
The discovery of the binary neutron star (BNS) merger GW170817 [1] marked a major milestone in multi-messenger astrophysics, as the observations of gravitational waves (GW) by LIGO-Virgo [2, 3] was followed by the detection of electromagnetic counterparts across multiple wavelengths [4, 5, 6, 7, 8]. These observations confirmed the connection between short gamma-ray bursts, kilonovae and BNS mergers and provided important information on the geometry of the jet launched after the merger as well as on the ejected neutron-rich material. Joint analysis of electromagnetic and GW data led to tight constraints of General Relativity [9], whereas the identification of the host galaxy allowed an independent measurement of the Hubble parameter (see, e.g., [10]). Combined with the second, GW-only, BNS detection [11], the local merger rate is constrained to $10-1700\,{\rm Gpc}^{-3}\,{\rm yr}^{-2}$ [12]. This information is crucial for building population models and understanding the formation of binary compact objects.
Observations of BNS mergers also provide a unique probe of the physics of extremely dense matter and allow to study the equations of state (EOS) of neutron star interiors. Important constraints on the EOSs can come from electromagnetic observations of isolated neutron stars or via GW observations of BNS mergers, as demonstrated with the detection GW170817. Since the EOS strongly affects the gravitational waveform, it leaves an imprint in the tidal deformation takes place in the late inspiral stages and allows measuring the tidal deformability [14, 15, 16, 17, 18]. In the case of GW170817, GW observations allowed to constrain the tidal deformability parameter to $\tilde{\Lambda}\lesssim 700$, where $\tilde{\Lambda}$ is the mass-weighted combination of the dimensionless deformability parameters [19, 20].
Additional and complementary information on the EOS can be extracted from the postmerger signal emitted by the hypermassive neutron star (HMNS) formed after the neutron stars have coalesced [21, 22, 23]. In particular, a number of recent studies have shown that the GW emission from the HMNS is concentrated at specific frequencies, resulting in relatively narrow spectral features [24, 25, 26, 27, 28, 29, 30]. The postmerger signal has not been detected so far, as the relevant frequencies are above those observable with current ground-based detectors. However, future third-generation (3G) detectors, such as the Einstein Telescope (ET) [31] and the Cosmic Explorer (CE) [CE: 32] will be able to detect and characterize the postmerger signal for sufficiently nearby systems.
A major limitation of GW observations of BNS mergers is that they are difficult to observe at large distances. Given typical BNS masses, current detectors are limited to only $\sim 500\,{\rm Mpc}$ [33]. While ET and CE will be able to reach $z\sim 2-3$ [34], in order to accurately measure the postmerger signal a high signal-to-noise ratio (SNR) is required, which is only achievable for relatively nearby systems (up to a few tens of Mpc) [35].
An alternative method to study populations of compact binaries is by observing the stochastic GW background (SGWB) created by the incoherent superposition of unresolved sources [36, 37]. Different types of sources can contribute to the SGWB, including compact binaries, core-collapse supernovae and isolated spinning neutron stars [e.g. 38, 39, 40]. Observing the SGWB will allow probing high-redshift populations, which are otherwise not accessible to ground-based detectors, although various methods are being developed to extract the SGWB from the detector noise in data from Advanced LIGO, Advanced Virgo and Kagra (LVK) are ongoing [41, 42]. Current upper limits for ground-based detectors are $\Omega_{\rm{GW}}(f)\leq 3.4\times 10^{-9}$ at $f=25$ Hz, where $\Omega_{\rm{GW}}$ is the dimensionless energy density of the SGWB [43, 44, 45, 46, 47]. These upper limits contain valuable information on high-redshift populations of compact binaries, complementary to that of individual detections [e.g. 48, 49, 50, 51, 52, 53, 54, 55]. For the LVK detectors, the dominant contribution to the SGWB is expected to be that of binary black holes. However, this will not be the case for ET and CE, since almost all the binary black hole mergers will be detected individually and subtracted from the data. The remaining SGWB will thus contain information on the unresolved BNS mergers.
In the case of the postmerger emission from a population of BNS mergers, we expect that if HMNS are formed, the characteristic GW frequencies will be enhanced in the SGWB due to the large number of unresolved sources contributing to the signal. This opens up the possibility to detect this EOS-dependent feature and to obtain complementary information on the postmerger stage. The goal of this study is to model the SGWB from a population of BNS, including the postmerger stage, and to evaluate its detectability with ET and CE. This paper is structured as follows. In Section II we present the BNS waveform used in this study, our method to construct a distribution of EOSs and discuss the calculation of the BNS background. We present our results in Section III and discuss future prospects in Section IV.
Figure 1: Evolution of the radiated GW energy when normalized to the at the merger. Shown with blue- and red-shaded areas are the corresponding values during the inspiral (blue) and the postmerger. The variance in the shaded regions marks the $99\%$ percentile and reflects the range of losses associated with the population of BNS considered here. Note that the postmerger energy can easily be $3-4$ larger than the inspiral one over the first $\sim 20\,{\rm ms}$.
II Methodology
II.1 BNS inspiral and postmerger waveform
A major focus of our work is to extend the contribution of the BNS merger signal to the SGWB by considering not only the inspiral part but also the postmerger emission. To appreciate simply why this is important and hence the relevance of our findings, we show in Fig. 1 the total GW energy radiated by a BNS when integrated over the entire frequency range ($E_{\mathrm{GW}}$) and normalized by the total energy considering only the inspiral phase ($E_{\mathrm{GW},0}$), as a function of the time before and after the merger $t-t_{\rm mer}$. Near the merger ($t-t_{\rm mer}\lesssim 0$) the energy radiated increases very rapidly, gaining an order of magnitude in the last $10\,{\rm ms}$ before the merger. Following the merger ($t-t_{\rm mer}>0$), the energy radiated continues to grow at a slower rate but more intensely, so that it can reach values much larger than those lost during the whole inspiral. Stated differently, BNS mergers are copious emitters of GWs during the inspiral, but it is the postmerger that provides the largest contribution to the radiated energy.
Of course, the exact amount of energy radiated will depend sensitively on the EOS describing the matter of which the BNS is composed of and which is essentially unknown. Instead of concentrating on a few specific EOSs, it is more convenient to consider a large ensemble of agnostic EOSs satisfy astrophysical, GW, or elementary-particle constraints [56, 57, 58, 59]. This is what is represented with the blue-shaded band in Fig. 1, which reports the 99th percentile contour of an ensemble of $10^{3}$ EOSs generated using our EOS-agnostic approach (see Sec. II.2). Similarly, the solid blue line reports the median of the distribution at each time in the evolution. In this way, it is straightforward to appreciate that since most of the energy is lost after the merger, it is important to include also the postmerger emission when calculating the BNS GW background (similar considerations apply when considering the radiated angular momentum [60, 61]).
To construct a hybrid waveform, we combine in the frequency domain the analytic prescriptions referring to the inspiral and the postmerger. As presented in detail in Appendix A, this combination takes place at a frequency $f_{\rm tr}$, where we move from the inspiral (see Sec. A.1) to the postmerger (see Sec. A.2) description. This frequency must be significantly smaller than the main frequencies appearing in the power spectral density (PSD) of the postmerger signal, i.e., $f_{1}$ and $f_{2}$, where $f_{1}<f_{2}$ (see Sec. A.2) so that no power from the postmerger phase is missed. Hereafter, for simplicity we set $f_{\rm tr}=f_{1}/2$, but the overall results depend only weakly on this assumption.
Figure 2 illustrates the construction of the hybrid waveform for a BNS system with masses $M_{1}=M_{2}=1.25\,M_{\odot}$, at a distance from observer $D=50\,{\rm Mpc}$ when employing GNH3 EOS discussed in Ref. [62]. Shown with a blue dashed line is the inspiral-and-merger signal followed by a prompt collapse and described by the IMRPhenomD_NRTidalv2 model, the postmerger signal with a black dashed line, and the combined (hybrid) signal with a red solid line. We also show with a solid green line the ET strain sensitivity. The transition frequency $f_{\rm tr}$ is meant to mark therefore the boundary between the grey and white regions. In the grey region, the hybrid waveform is simply given by the IMRPhenomD_NRTidalv2 waveform. Note that in the white region, the hybrid waveform is given by the sum of the inspiral and postmerger strains and that the latter is negligible below $f_{\rm tr}$, thus ensuring no significant power is lost or added. To properly combine the two waveforms above $f_{\rm tr}$, we must ensure phase coherence, which is not guaranteed a priori, as we have not explicitly computed the temporal phase $\beta_{2}$ in Eq. (8). Therefore, at $f_{\rm tr}$, we adjust the phase of the IMRPhenomD_NRTidalv2 waveform to match the postmerger one, thus ensuring phase continuity.
Figure 2: Evolution in the frequency-domain of the GW strain for a BNS with masses $m_{1}=m_{2}=1.25\,M_{\odot}$ at a distance of observer $D=50\,{\rm Mpc}$ and relative to the GNH3 EOS [62]. The IMRPhenomD_NRTidalv2 waveform is shown with a black dashed line, the HMNS postmerger with a blue dashed line, and the combined waveform with a red solid line. The frequency $f_{\rm tr}$ marks the transition between the inspiral (grey-shaded are) and the postmerger, while the green solid curve reports the sensitivity of ET
II.2 Modelling the equations of state
In view of the uncertainties in the EOS and their effect on the waveform, we adopt an agnostic approach in this work. We use the findings of Refs. [59, 63, 64, 65], who constructed more than $10^{7}$ EOSs consistent with constraints from fundamental physics and astronomical observations of neutron stars. Leveraging on this extensive dataset, it was possible to derive informative constraints on the relationship between the chirp mass $\mathcal{M}_{\mathrm{chirp}}$ and tidal deformability $\Lambda$ in terms of a probability density function (PDF) for these two quantities. We have exploited this correlation to create a consistent and agnostic description, assigning a range of allowed values of the tidal deformability for any choice of the chirp mass.
A-priori, a single EOS does not correspond to a unique line in the $\mathcal{M}_{\mathrm{chirp}}$-$\Lambda$ plane because the same chirp mass may result from many combinations of BNS masses. However, when restricting the analysis to symmetric binaries, as we do throughout the rest of this work, this ambiguity disappears. In this case, for a given EOS, each chirp mass is uniquely associated to a single value of $\Lambda$. Therefore, extracting an EOS from a given distribution is equivalent to drawing a single value of $\Lambda$ from the PDF at each chirp mass. Similarly, extracting a certain value of $\Lambda$ from the allowed range is equivalent to singling-out a specific EOS.
Although the PDF found in Ref. [59] for a given chirp mass is not exactly a Gaussian, it is so to a first approximation [see Fig. (4) in [59]]. Hence, when assuming an underlining Gaussian distribution, we can relate the tidal deformability and the chirp mass (in solar masses) via a simple power-law of the type
$\Lambda_{p}=\alpha+\beta\mathcal{M}^{\gamma}_{\mathrm{chirp}}\,,$ (1)
where $\Lambda_{p}$ is normally distributed between the minimum and maximum values given by [59]
$\Lambda_{\rm min\,(max)}=a+b\,\mathcal{M}_{\rm chirp}^{c}\,,$ (2)
with $a=-50\,(-20)$, $b=500\,(1800)$, and $c=-4.5\,(-5.0)$. In this way, we can build a set of fictitious $10^{3}$ EOSs such that the corresponding tidal deformability follows expression (1) and is normally distributed.
Once an EOS is set, we can compute the inspiral waveform as discussed in Sec. A.1. For the postmerger waveform, and as discussed in Sec. A.2, we need to know the characteristic frequencies $f_{1}$ and $f_{2}$ given the selected EOS and hence tidal deformability. Luckily, this is rather straightforward to do since universal relations have been found that provide analytic fits (see, e.g., [25, 26, 27, 28]). Here, we employ the expressions proposed in Refs. [28, 30] (and also employed in [62]) and express the postmerger frequencies $f_{1}$ and $f_{2}$ as
$\frac{f_{1}}{\mathrm{kHz}}$ $=$ $c_{0}+c_{1}x+c_{2}x^{2}+c_{3}x^{3}\,,$ (3)
$\frac{f_{2}}{\mathrm{kHz}}$ $=$ $10^{\,a_{0}+b_{0}\,x}\,\left(\frac{M_{\odot}}{M}\right)\,,$ (4)
where $x:=(3\Lambda/16)^{1/5}$, $M:=m_{1}+m_{2}$ is the total mass of the system, and the constants are $c_{0}=45.195$, $c_{1}=-43.484$, $c_{2}=14.653$, $c_{3}=-1.662$, $a_{0}=1.617$, and $b_{0}=-0.274$.
We conclude this section by remarking that expressions (4), being only “quasi”-universal relations, obviously come with some intrinsic uncertainties as estimated from the numerical simulations and the corresponding fits (see discussion in [30]). These uncertainties, however, do not play an important role here as we are interested in a SGWB that is naturally broad given the much larger uncertainties in the EOS.
II.3 Stochastic gravitational-wave background from BNS mergers
The energy density in GWs per logarithmic unit of frequency and in units of the critical density of the Universe, $\rho_{\mathrm{c}}=3H_{0}/8\pi G$, is given by
$\Omega_{\mathrm{GW}}:=\frac{1}{\rho_{c}}\frac{\mathrm{d}\rho_{\mathrm{GW}}}{\mathrm{d}\ln{f}}=\frac{f}{c\,\rho_{\mathrm{c}}}F(f)\,,$ (5)
where $\rho_{\mathrm{GW}}$ is the GW energy density and $F(f)$ is the energy flux. When considering a population of $N$ sources, the energy flux per unit frequency is given by [66]:
$F(f):=\frac{\pi c^{3}}{2G}\frac{f^{2}}{T}\sum_{i=1}^{N}\left[|\tilde{h}_{i}^{+}(f)|^{2}+|\tilde{h}_{i}^{\times}(f)|^{2}\right]\,,$ (6)
where the index $i$ sums over the $N$ sources and $\tilde{h}_{i}^{+}(f)$ and $\tilde{h}_{i}^{\times}(f)$ are the Fourier transforms of the two GW polarization modes. The number of sources is taken to be proportional to the total observation time $T$, which in this study is set to one year. We have checked that our calculation converges at $T=1$ year and will use this value in the following (note that the background is stationary).
In order to construct our BNS catalogue, we follow the baseline_delays model from Ref. [55] and, more specifically, we assume all binaries are symmetric in mass and follow a Gaussian mass distribution with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. This distribution is consistent with the theoretical upper limits on the maximum mass of a nonrotating NS of $\sim 2.16-2.3M_{\odot}$, in particular the constraints obtained from multi-messenger observations of GW170817 [67, 68, 69, 70, 71].
The BNS-merger rate is described by the star-formation rate from Ref. [72] convolved with a distribution of delay times $t_{\mathrm{d}}$ that follows a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ [73]. Using this model, we construct a catalogue containing about $N=3.8\times~{}10^{5}$ BNS that merge in the entire Universe during one year, consistent with other studies [74] (see Appendix B and Ref. [55]).
As we focus on the background, we must consider the flux contributions from non-detectable sources only. The detectability of a GW signal is estimated by calculating its SNR that, for a single GW detector, is given by
$\mathrm{SNR}^{2}:=4\int^{\infty}_{0}\frac{|\tilde{h}(f)|^{2}}{P_{\mathrm{n}}(f)}\,\mathrm{d}f\,,$ (7)
where $P_{\mathrm{n}}(f)$ is the noise PSD of the detector. For a network of $N_{\mathrm{det}}$ detectors, and assuming an uncorrelated noise among them, the SNR of a GW signal is calculated by summing the individual-detector SNRs in quadrature; we have here employed the GWFish package [75] to compute the SNR.
We assume that individual resolvable sources can be identified when their SNR exceeds a threshold $\mathrm{SNR}_{\mathrm{thr}}$, so that the contribution from these resolvable sources is removed by subtracting each individual signal, however the process of parameter estimation of these sources inevitably comes with an uncertainty, so that the subtraction process is inherently affected and may introduce residual contributions to the background, as discussed in Ref. [76]. In this study, which represents a first investigation of the postmerger contribution, we assume that the subtraction is noiseless, and leave more realistic approaches for future work [77, 78, 79].
Following Refs. [80, 74], we take as SNR threshold for the detection of BNS mergers $\mathrm{SNR}_{\mathrm{thr}}=12$, but remark that for 3G detectors this value will depend on the details of the analysis pipeline and may affect the efficiency of signal subtraction. Indeed, it will be difficult to accurately measure the parameters of low-SNR sources, whose subtraction will introduce more noise in the SGWB compared to high-SNR sources. In this work, we investigate the following values for $\mathrm{SNR}_{\mathrm{thr}}=12,20,40$, and $80$.
III Results
Figure 3 shows the SGWB from BNS mergers computed with our two descriptions for the modelling of the signal and detailed in Sec. II.1. More specifically, in the inspiral+prompt collapse model, hereafter “IPC” (blue-shaded areas), which is the one adopted so far in describing the SGWB from BNS mergers, the inspiral is assumed to be followed by a prompt collapse to a BH of the merger remnant. On the other hand, in the novel inspiral+postmerger model, hereafter “IPM”, proposed here (red-shaded areas), the signal includes the postmerger emission from a metastable HMNS that eventually collapses into a BH. For both models, we use ${\rm SNR}_{\mathrm{thr}}=12$, and the background was calculated using $10^{3}$ randomly sampled EOSs using the methodology described in Sec. II.2. The shaded regions correspond to the $68$ and $95$ percentiles of the resulting distribution, respectively, while the solid lines indicate the median values in the distributions.
Figure 3: Energy density of the SGWB from BNS mergers, corresponding to either the IPC (blue-shaded areas) or the IPM (red-shaded areas) scenarios considered here. Shown with different transparencies are the $68$ and $95$ percentile contours, while the solid lines report the medians. Also reported are the power-integrated sensitivity curves of a single ET (dashed lines) and of a network composed of an ET detector and of two CE detectors (solid lines) relative to $1$ (grey lines) and $10$ (black lines) years of integration, respectively.
Figure 4: Left panel: Energy density of the SGWB from BNS mergers in the IPM scenario. Shown with different transparencies are the $68,95$, and $99$ percentile contours, while the solid lines report the medians. Different colours are used to show how $\Omega_{\rm GW}$ changes when assuming threshold SNR values of ${\rm SNR_{thr}}=12,20,40$, and $80$; note that $\tau_{2}=50\,{\rm ms}$ in all cases. Also reported are the power-integrated sensitivity curves of a single ET (dashed lines) and of a network composed of an ET detector and of two CE detectors (solid lines) relative to $1$ (grey lines) and $10$ (black lines) years of integration, respectively. Right panel: The same as in the left but when considering different values for the decay timescale $\tau_{2}=20,50$ and $250\,{\rm ms}$ (blue, green, and red, respectively). Note that ${\rm SNR_{thr}}=12$ in all cases.
As shown in Fig. 3, both models coincide at low frequencies (i.e., below $200\,{\rm Hz}$), which is expected since their inspiral parts are identical. They begin to diverge at approximately $350\,{\rm Hz}$. The power in the IPC model rapidly decreases as expected, since it does not include significant postmerger emission. In contrast, the IPM model exhibits a peak followed by a gradually declining plateau, and then experiences a rapid rise to a broad peak near $1.7\,{\rm kHz}$, which is about six times higher than the $350\,{\rm Hz}$ peak in $99\%$ of the EOSs considered. This peak is broad because of the redshift distribution of the sources and the intrinsic differences in the $10^{3}$ EOSs considered. We note that the high-frequency peak arises solely from the postmerger contribution, while the low-frequency peak is primarily influenced by the inspiral phase, as indicated by the overlap of the blue and red distributions. On the other hand, the width of the SGWB peak is due mostly to the redshift distribution of the sources, while the uncertainty in the EOSs contained in our sample is what leads to the variance shown with the shaded areas.
Black solid and dashed lines in Fig. 3 represent the power-law integrated sensitivity curves of two 3G detector configurations (see Appendix C), namely, a single ET configuration (dashed lines) and an admittedly optimistic configuration consisting of two ET detectors, as well as two CE detectors (solid line), such type of network can drastically improve the background sensitivity [81, 82, 83]. Sensitivity curves for both scenarios are shown either for one-year integration time (grey solid lines) or for ten years of observations (black lines) and with an SNR for SGWB detection of $\rho=1$ [see Eq. (13)].
The results in Fig. 3 clearly indicate that even after ten years, a single 3G detector will not be able to reach the peaks in the SGWB arising from the inspiral or from the postmerger. However, the network setup could detect the low-frequency peak with just one year of data and potentially the high-frequency peak with ten years of observation. Note that the improvement in sensitivity between the two scenarios is less pronounced than expected, primarily because ET and CE are geographically distant, leading to a suppressed overlap-reduction function at high frequencies, thereby limiting the potential gains.
We have so far assumed that all the individually detected sources are subtracted noiselessly from the data, leaving out the clear SGWB from unresolved BNS. In practice, however, this subtraction is challenging, especially for low-SNR sources, whose parameters are recovered with large uncertainties. To assess the impact that this subtraction has on the estimated SGWB, we consider scenarios in which the individual detection threshold is increased, which is equivalent to tagging as detected only the loudest sources. In this case, we expect better source subtraction, leading to less noise in the SGWB, which will also contain more sources.
We explore different detection thresholds in the left panel of Fig. 4, where we compare our default value $\mathrm{SNR}_{\mathrm{thr}}=12$ with higher values of $20,40$, and $80$. As expected, the amplitude of the SGWB increases with $\mathrm{SNR}_{\mathrm{thr}}$, which is encouraging as the background approaches the network sensitivity. Indeed, as we increase $\mathrm{SNR}_{\mathrm{thr}}$, fewer sources are identified as individual detections, and instead contribute to the background. At the same time, as $\mathrm{SNR}_{\mathrm{thr}}$ is increased, the positions of both the low- and high-frequency peaks shift to higher frequencies. This shift occurs because the background increasingly includes brighter sources, which are typically closer and less redshifted. Therefore, the peaks move to larger frequencies, where the detector sensitivity is lower, so that the actual detection of the SGWB at these frequencies will remain challenging despite the decrease in the background noise.
[…]
As expected, the PSD in the high-frequency peak of the postmerger signal increases linearly with the increase of $\tau_{2}$, thus considerably increasing the potential of detection of this part of the SGWB. At the same time, integration times longer than one year are needed to detect the high-frequency portion of the SGWB related to the postmerger. Interestingly, however, the increase in the decay time also leads to an increased power at lower frequencies, i.e., $500\lesssim f\lesssim 1000\,{\rm Hz}$, so that the detectability of the high-frequency portion of the SGWB is possible even if the maximum of the postmerger PSD is still below sensitivity. In turn, the results shown in the right panel of Fig. 4 also reveal that the detection of the high-frequency portion of the SGWB can provide precious information on the lifetime of the remnant and hence on the EOS. It is unlikely that stringent constraints on the EOS can be drawn from this measurement, which however will serve as an important consistency check. We will explore the implications of these results and how they can be used to infer the nuclear-matter EOS in a future work.
IV Conclusion
The SGWB generated by the inspiral and merger of binary neutron stars has so far been modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for very massive binaries, it is not what is expected to take place for systems having masses around the peak of the mass distribution of neutron-star binaries. Hence, the most likely outcome of the merger is that of producing a remnant that will survive for hundreds of milliseconds before collapse (in the case of the GW170817 event, the merger lifetime has been estimated to be of about one second [84, 85]). During its short existence, the remnant can radiate enormous amounts of energy in GWs and, indeed, the radiated GW energy is 3-4 times larger than that emitted over the long inspiral.
These considerations have motivated us to include also the postmerger contribution to the PSD, and hence calculate how the SGWB is modified by this important and dominant contribution. To this scope, we have implemented an analytic model of the postmerger GW signal that has been shown to provide a reasonably accurate description of the signal computed by numerical-relativity simulations [62]. Over the timescales considered here, this analytic waveform is dominated by the fundamental frequency of oscillation of the remnant $f_{2}/2$, that provides a narrow-banded emission around $f_{2}\simeq 1-2\,{\rm kHz}$. Quasi-universal relations have then been employed to relate the fundamental frequency with the tidal deformability extracted from a large sample of $10^{3}$ EOSs built in a parametric way and satisfying all the known astrophysical and microphysical constraints. Additional quasi-universal relations have also been employed to relate the tidal deformability to the chirp mass of the binary. Finally, the analytic description of the GW emission and the parametric construction of the EOSs have then been combined with a model for the population of equal-mass BNS binaries [55]. In this way, the postmerger contribution depends mostly on only two parameters: the decay timescale of the GW at the $f_{2}$ frequency, i.e., $\tau_{2}$, and the threshold value of the SNR at which a BNS can be resolved out of the background, i.e., ${\rm SNR_{thr}}$. Because both $\tau_{2}$ and ${\rm SNR_{thr}}$ are poorly known, in our analysis they are varied across the reasonable intervals in which they are expected to range.
Using this setup, we were able to compute the high-frequency postmerger contribution to the SGWB from about $\sim 3.8\times 10^{5}$ binaries (i.e., those that are expected to merge in the whole universe in one year [55]), finding that this indeed represents a new and dominant contribution. While being broad, this contribution peaks at $\sim 1-2\,{\rm kHz}$ and provides a contribution to the normalised GW energy density $\Omega_{\rm{GW}}\simeq 10^{-10}-10^{-9}$, hence much larger than what is expected when considering scenarios with the inspiral and collapse only. Clearly, being a background signal, the detectability of this contribution will also depend on the integration time employed and our analysis has shown that long integration times and a network of 3G detectors are likely to be needed for the detection. However, such a network setup could detect the peak of the high-frequency component in less than one year if $\tau_{2}\gtrsim 100\,{\rm ms}$. More importantly, the detection of the high-frequency portion of the SGWB can provide precious information on the lifetime of the remnant and hence on the EOS. Overall, the results presented in this first analysis show that the postmerger contribution to the SGWB cannot be ignored as done so far, and that it contains crucial information on the HMNS. It is therefore important to extend the modelling of the SGWB to higher frequencies and to consider the detection of this background with 3G observatories.
Notwithstanding the innovative nature of the procedure presented here, our approach can be improved in a number of ways. First, by considering that the decay time is not a simple mass dependant power-law, but actually a function of the EOS. Second, by extending the analysis to binaries that are not symmetric and hence introduce a degeneracy in the modelling of the chirp mass. Finally, by considering a more careful analysis of the foreground subtraction, which has been here treated simply in terms of the threshold SNR. While we do not expect any of these improvements to change the qualitative aspects of our results, they could impact the quantitative ones. We will explore these refinements of the model in future work.
V Acknowledgements
This research is supported by the École Normale Supérieure Paris-Saclay CDSN PhD grant. This research is also supported in part by the ERC Advanced Grant “JETSET: Launching, propagation and emission of relativistic jets from binary mergers and across mass scales” (grant No. 884631), by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 “Strong-interaction matter under extreme conditions” – project number 315477589 – TRR 211, and by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). LR acknowledges the Walter Greiner Gesellschaft zur Förderung der physikalischen Grundlagenforschung e.V. through the Carl W. Fueck Laureatus Chair.
Appendix A Modelling the BNS waveforms
A.1 Modelling the Inspiral
To model the inspiral part of the GW waveform, we use the IMRPhenomD_NRTidalv2 model [17, 18] from the LALSuite simulation package [86]. This model extends the IMRPhenomD waveform [87, 88], designed for binary black hole systems, by incorporating EOS-dependent tidal effects. Specifically, the template uses the dimensionless tidal deformability parameter $\Lambda:=G^{4}\lambda/c^{5}M^{5}$, where $M$ is the stellar mass and the tidal deformability parameter $\lambda$ is defined as the ratio between the induced quadrupole and the external gravitational field ($G$ and $c$ are the gravitational constant and the speed of light, respectively). The model assumes that the merger remnant promptly collapses to a black hole, thus leading to very little power in the postmerger phase [18]. As commented above, this is a very drastic assumption and applicable only to a rather restricted set of binaries with total gravitational mass that is roughly given by $M_{\rm tot}\gtrsim 1.45\,M_{{}_{\rm TOV}}\,\sim 3.38\,M_{\odot}$ [89], where $M_{{}_{\rm TOV}}$ is the maximum nonrotating gravitational mass for a given EOS and the estimate comes from assuming a maximum mass $M_{{}_{\rm TOV}}\sim 2.33\,M_{\odot}$ [67, 68] (see also [90, 91, 92, 93] for a discussion on the threshold mass in BNS mergers).
A.2 Modelling the Postmerger
Numerical-relativity simulations suggest that BNS mergers typically produce a metastable HMNS, which persists for several tens of milliseconds before ultimately collapsing into a spinning black hole [22, 94]. Recent studies showed that the GW signal emitted by this HMNS exhibits several prominent peaks in the frequency domain [24, 25, 27, 28, 26, 29, 30].
Here, we use the results of Ref. [62], which presents analytical fits for the postmerger GW signal based on an extensive catalogue of time-domain numerical waveforms. These fits express the time-domain strain as a sum of damped sinusoids with a time-evolving instantaneous frequency as follows:
$h_{+}(t)=h_{0}\,\mathrm{e}^{-t/\tau_{1}}\left[\mathrm{sin}(2\pi f_{1}t)+\mathrm{sin}(2\pi(f_{1}-f_{1\epsilon})t)\right.\,\\ \hskip 59.46614pt\left.+\mathrm{sin}(2\pi(f_{1}+f_{1\epsilon})t)\right]\\ +\mathrm{e}^{-t/\tau_{2}}\,\mathrm{sin}\left(2\pi f_{2}t+2\pi\gamma_{2}t^{2}+2\pi\xi_{2}t^{3}+\pi\beta_{2}\right)\,,$ (8)
where $h_{0}=G\mathcal{M}_{\mathrm{chirp}}/(2Dc^{2})$ with $D$ the distance to the source, $t=0$ is the merger time, the frequencies $f_{1}$ and $f_{2}$ represent the dominant postmerger GW modes with $f_{2}>f_{1}$, $\tau_{1}$ and $\tau_{2}$ are the decay times of each mode, and $f_{1\epsilon}$, $\beta_{2}$, $\gamma_{2}$ and $\xi_{2}$ are constants [24, 95, 28]. The first peak, $f_{1}$, is short-lived (with a decay time $\tau_{1}$ of the order of a few ms) and most probably originates from damped collisions of the two stellar cores immediately after the merger [96]. This peak is accompanied by a symmetric peak $f_{3}\sim 2f_{2}-f_{1}$, which we do not take into account here, as its frequency is very high and unlikely to be detected even by third-generation (3G) detectors. The second peak, $f_{2}$, corresponds to the spin frequency of the deformed HMNS and has a decay time $\tau_{2}$ of the order of several tens of milliseconds [95, 97]. However, the expected values, and mass dependencies of both $\tau_{1}$ and $\tau_{2}$ are still rather uncertain as they sensitively depend on the EOS and the microphysical effects taking place after the merger. Interestingly, when considering the information collected with the GW170817 event, some studies have constrained the maximum lifetime of the remnant from that event to approximately one second [84, 85]. Overall, it is clear that the lifetime of the HMNS can vary significantly between a few to several thousands of milliseconds.
Here, we parametrize both decay times as power laws of the form
[…]
where $\mathcal{M}_{\mathrm{chirp}}:=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass and $m_{1}$, $m_{2}$ are the component masses. Note that $\gamma<0$, as higher chirp masses correspond to a more rapid collapse, and reference decay times are taken to be $\tau_{1}=5\,{\rm ms}$ and $\tau_{2}=50\,{\rm ms}$ which correspond to typical lifetimes values Ref. [62]. Because a representative value of $\tau_{2}$ is only poorly known and the value adopted above is quite conservative, we will consider later on how our results are affected when considering larger values of $\tau_{2}$. Finally, as a reference in our analysis we take the value inferred for GW170817, i.e., $\mathcal{M}_{\mathrm{chirp,ref}}=1.186\,M_{\odot}$, and set $\gamma=-3$; alternative parametrizations are discussed in Section III, where we evaluate the impact of varying $\tau_{2}$.
The parameter $f_{1\epsilon}=50\,{\rm Hz}$ introduces a small frequency modulation in the $f_{1}$ mode, while $\beta_{2}$ adjusts the temporal phase of the waveform to ensure smooth continuity between inspiral and postmerger phases. In this work we set $\beta_{2}=0$ for simplicity and adjust the phase in the frequency domain as explained below. Furthermore, higher-order contributions to the waveform modelled by Eq. (8) come from the coefficients $\gamma_{2}$ and $\xi_{2}$, but these have a relatively minor impact on the total GW power. We take these parameters to be constant and equal to their mean values over the four EOSs studied by Ref. [62], thus $\gamma_{2}=-554\,{\rm Hz}^{2}$ and $\xi_{2}=3.36\times 10^{4}\,{\rm Hz}^{3}$.
Appendix B Modelling the BNS population
In order to calculate the SGWB from the entire population of unresolved BNS mergers, we need to model the merger rate as a function of redshift and system parameters. Here, we consider (equal-mass) symmetric non-eccentric binaries, so that each BNS is characterized by the component masses $m_{1}=m_{2}=m$, tidal deformability parameters $\Lambda_{1}=\Lambda_{2}=\Lambda$, redshift $z$, sky position (right ascension $\alpha$ and declination $\delta$), orbital inclination angle $\iota$, polarization angle ($\psi$), coalescence phase $\Phi_{\mathrm{c}}$ and coalescence time set to be $t_{\mathrm{c}}=0$. The mass distribution we used is derived from Galactic observations and is assumed to be independent of redshift. According to Ref. [98], this distribution is well-modeled by a Gaussian with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. Both neutron stars in the binary are assumed to follow this distribution, which is tightly concentrated around the mean, making it a non-critical parameter for this study.
We note that this mass distribution may depend on the formation scenario of the binary. More specifically, the mass distribution of BNS detected via GW may differ from that of Galactic binaries, potentially explaining the higher total mass observed in the BNS merger GW190425 [11, 99]. Furthermore, the recent GW190814 event suggests a black hole-neutron star (BHNS) merger involving a neutron star significantly heavier than typical Galactic counterparts [100], although there are arguments favouring a black-hole origin for the secondary [71].
We assume that the BNS merger rate follows the star formation rate (SFR) convolved with a distribution of delay times. Specifically, we adopt the baseline_delays model from Ref. [55], where the merger rate is defined as:
$R_{\mathrm{merg}}(t)=R_{0}\int^{t_{\mathrm{d},\mathrm{max}}}_{t_{\mathrm{d},\mathrm{min}}}\mathrm{SFR}(t-t_{\mathrm{d}})\,P(t_{\mathrm{d}})\,\mathrm{d}t_{\mathrm{d}}\,.$ (10)
Here the SFR is taken from Ref. [72] and $P(t_{\mathrm{d}})$ is the distribution of time intervals $t_{\mathrm{d}}$ between the formation of stellar progenitors and the eventual BNS merger. We here assume a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ [73], and the delay times to range from $t_{\mathrm{d},\mathrm{min}}=10\,{\rm Myr}$ to $t_{\mathrm{d},\mathrm{max}}$ set by the Hubble time. The parameter $R_{0}$ in Eq. (10) is a normalization constant that calibrates the merger rate at $z=0$ the local BNS merger rate inferred from the GWTC-3 LIGO-Virgo catalogue: $R_{0}=44^{+96}_{-34}~{}~{}{\rm yr}^{-1}\,{\rm Gpc}^{-3}$ [12].
Appendix C Network sensitivity for SGWB
The SGWB is in noise-dominated mode for ground-based detectors, as it is always much weaker than the detector noise. Moreover, its stochastic nature makes the use of match-filtering techniques, which have been developed and are effective for individual sources, inadequate. Instead, detection methods rely on cross-correlations of the outputs of pairs of detectors and searches for excess signal, under the assumption that the noise is only weakly correlated [101, 102]. Knowledge of the expected signal spectrum can help in constructing adapted searches, and if the spectrum is broad, integration over frequencies can increase the SNR. In order to better visualise the detectability of the SGWB, we use power-integrated (PI) sensitivity curves, following Ref. [103] and briefly summarised below. We also outline our calculation of PI sensitivity curves for a network of ET and CE detectors with different configurations.
We start by assuming the background has a power-law spectrum with slope $\sigma$ and amplitude $\Omega_{0}$
$\Omega_{\rm{GW}}(f)=\Omega_{0}\left(\frac{f}{f_{\rm{ref}}}\right)^{\alpha}\,.$ (11)
It can then be shown that $\alpha=2/3$ for the inspiral part of the spectrum. We further define the PSD $S_{h}(f)$ of the GW signal as [66]
$S_{h}(f):=\frac{3H_{0}^{2}}{2\pi^{2}}\frac{\Omega_{\rm{GW}}(f)}{f^{3}}\,,$ (12)
where $H_{0}$ is the Hubble parameter, and we are using a single-sided spectrum.
The broadband SNR for a network cross-correlation search for an unplaced and isotropic stochastic background $\Omega_{\mathrm{GW}}(f)$, integrated over frequency in the range $[f_{\mathrm{min}},f_{\mathrm{max}}]$, is given by [103]
$\rho:=\sqrt{2\tau}\left[\int^{f_{\mathrm{max}}}_{f_{\mathrm{min}}}\sum_{i=1}^{N}\sum_{j>i}^{N}\frac{\gamma^{2}_{ij}(f)S_{h}^{2}(f)\sin^{2}\beta_{i}\sin^{2}\beta_{j}}{P_{n,i}(f)P_{n,j}(f)}\mathrm{d}f\right]\,,$ (13)
where $\tau$ is the observation time, $\beta_{i}$ and $\beta_{j}$ are the opening angles of interferometers $i$ and $j$, respectively, $P_{n,i}$ and $P_{n,j}$ are their noise PSDs, and $\gamma(f)$ is the overlap reduction function that quantifies the correlation between the signals observed by the different interferometers [104].
It is important to note that for detectors that are located far apart or have large relative orientation angles, $\gamma(f)$ becomes an oscillatory function that approaches zero at high frequencies. This behaviour reduces the sensitivity of the detector pair in the frequency range of interest for this study (i.e., above $\sim 1\,{\rm kHz}$). Conversely, for co-located detectors, $\gamma(f)$ remains constant across all frequencies and varies only with the orientation angle, reaching a maximum value of $\gamma=1$ for co-aligned L-shaped detectors.
The PI sensitivity curve, constructed for a given SNR $\rho_{0}$ and observation time $\tau$, is defined such that a power-law SGWB tangent to this curve is detected with this SNR for an observation time $\tau$. More specifically, for a given slope $\alpha$ of the SGWB and for a given observation time $\tau$, we can find an amplitude $\Omega_{0,\alpha}$ such that the SNR in Eq. (13) is given by some fixed value $\rho=\rho_{0}$
$\Omega_{0,\alpha}:=\frac{\rho_{0}}{\sqrt{2\tau}}\,\left[\int^{f_{\mathrm{max}}}_{f_{\mathrm{min}}}\frac{(f/f_{\mathrm{ref}})^{2\alpha}}{\Omega_{\mathrm{eff}}^{2}}\mathrm{d}f\right]^{-1/2}\,.$ (14)
lat () lon () az () $\beta$ () $L({\rm km})$
${\rm ET}$ $40.5$ 9.49.4~{}~{}~{}9.49.4 $70.6$ $60$ $10$
${\rm CE}_{1}$ $46.5$ $-119.4$ 126126126~{}126 $90$ $40$
${\rm CE}_{2}$ $46.5$ $-119.4$ 126126126~{}126 $90$ $20$
Table 1: Latitude, longitude, azimuth, opening angle and arms length of the interferometers considered in this study. The corresponding noise PSDs are taken from Refs. [31, 105].
The effective background $\Omega_{\mathrm{eff}}$ is defined using Eq. (12) and the effective-strain noise PSD $S_{\mathrm{eff}}$ as [81]
$S_{\mathrm{eff}}(f):=\left[\sum_{i=1}^{N}\sum_{j>i}^{N}\frac{\gamma^{2}_{ij}(f)\sin^{2}\beta_{i}\sin^{2}\beta_{j}}{P_{n,i}(f)P_{n,j}(f)}\right]^{-1/2}\,,$ (15)
where, for a single interferometer, Eq. (15) reduces to $S_{\mathrm{eff}}=\sqrt{P_{\mathrm{n}}}=S_{\mathrm{n}}$, as expected. The PI sensitivity curve is then given by the envelope of all the power-law functions $\Omega_{\rm{GW}}$ with $\Omega_{0,\alpha}$ and $\alpha$ related through Eq. (14). We refer the reader to Ref. [103] for further details on this method. In the following, all the sensitivity curves we report are PI sensitivity curves.
Two different scenarios have been considered in this work: a single triangular $10\,{\rm km}$ ET detector and a network composed of 2 ET detectors ${\rm ET}_{1,2}$ combined with two CE detectors ${\rm CE}_{1,2}$. The latter case includes two triangular $10\,{\rm km}$ ET and two CEs: a $20\,{\rm km}$ L-shaped ${\rm CE}_{1}$ and a $40\,{\rm km}$ L-shaped ${\rm CE}_{2}$. All the parameters for each scenario are detailed in Tab. 1, while the PSDs for ET and CE are taken from Refs. [31] and [105], respectively. We note that we neglect the impact of correlated noise, which may however be significant for co-located ET detectors [31, 106]. Finally, we optimistically assume the detectors to operate at a duty cycle of $100\%$, and explore two possibilities for the observations time: $1$ and $10$ years.