### v2

# Postmerger: a new and dominant contribution to the gravitational-wave background from binary neutron stars

Abstract
The stochastic gravitational-wave background (SGWB) generated by the inspiral and merger of binary neutron stars is traditionally modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for the most massive binaries, it is not what is expected in general, as the remnant may survive for up to hundreds of milliseconds and radiate an amount of energy that is significantly larger than that lost during the whole inspiral. To account for this additional contribution to the SGWB, we consider a waveform model that includes both the inspiral and the postmerger emission. We show for the first time that for a large set of parameterized equations of state (EOSs) compatible with observational constraints, there is considerable spectral power in the $1-2\,{\rm kHz}$ range, distinct from that associated with the inspiral and leading to a dimensionless GW energy density $\Omega_{\rm GW}\simeq 10^{-10}-10^{-9}$. We discuss the enhanced detectability of the SGWB by third-generation detectors such as the Einstein Telescope and Cosmic Explorer, and show how it depends on the signal-to-noise ratio of foreground binaries and on the remnant lifetime. Interestingly, even a non-detection of the high-frequency part of the SGWB could provide valuable constraints on the remnant lifetime, offering novel insights into the postmerger dynamics and the EOS of nuclear matter.
Introduction— The discovery of the binary neutron star (BNS) merger GW170817 LIGO and Virgo collaboration (2017a) marked a major milestone in multi-messenger astrophysics, as the observations of gravitational waves (GW) by LIGO-Virgo LIGO collaboration (2015); Acernese, F. et al. (2015) was followed by the detection of electromagnetic counterparts across multiple wavelengths Coulter, D. A. et al. (2017); Goldstein, A. et al. (2017); Savchenko, V. et al. (2017); Mooley et al. (2018); Ghirlanda, G. et al. (2019). These observations confirmed the connection between short gamma-ray bursts, kilonovae and BNS mergers and provided important information on the geometry of the jet launched after the merger as well as on the ejected neutron-rich material. Joint analysis of electromagnetic and GW data led to tight constraints of General Relativity collaboration (2019), whereas the identification of the host galaxy allowed an independent measurement of the Hubble parameter (see, e.g., LIGO and Virgo collaboration (2017b)). Combined with the second, GW-only, BNS detection collaboration (2020), the local merger rate is constrained to $10-1700\,{\rm Gpc}^{-3}\,{\rm yr}^{-2}$ Abbott, R. et al. (2023). This information is crucial for building population models and understanding the formation of binary compact objects.
Observations of BNS mergers also provide a unique probe of the physics of extremely dense matter and allow to study the equations of state (EOS) of neutron star interiors. Since the EOS strongly affects the gravitational waveform, it leaves an imprint in the tidal deformation that takes place in the late inspiral stages and allows measuring the tidal deformability Bernuzzi et al. (2015); Hinderer, T. et al. (2016); Radice et al. (2014); Dietrich, Tim et al. (2019); Dietrich et al. (2019). In the case of the GW170817 event, GW observations allowed to constrain the tidal deformability parameter to $\tilde{\Lambda}\lesssim 700$, where $\tilde{\Lambda}$ is the mass-weighted combination of the dimensionless deformability parameters Abbott, B. P. et al. (2018a); LIGO and Virgo collaboration (2019).
Additional and complementary information on the EOS can be extracted from the postmerger signal emitted by the hypermassive neutron star (HMNS) formed after the neutron stars have coalesced Shibata et al. (2017); Baiotti and Rezzolla (2017); Radice et al. (2020). In particular, a number of recent studies have shown that the GW emission from the HMNS is concentrated at specific frequencies, resulting in relatively narrow spectral features Stergioulas et al. (2011); Bauswein and Janka (2012); Takami et al. (2014a); Bauswein and Stergioulas (2015a); Rezzolla and Takami (2016); Takami et al. (2015a); Topolski et al. (2024). The postmerger signal has not been detected so far, as the relevant frequencies are above those observable with current ground-based detectors. However, future third-generation (3G) detectors, such as the Einstein Telescope (ET) Maggiore, M. et al. (2020) and the Cosmic Explorer (CE) Evans, M. et al. (2021) will be able to detect and characterize the postmerger signal for sufficiently nearby systems.
An alternative method to study populations of compact binaries is by observing the stochastic GW background (SGWB) created by the incoherent superposition of unresolved sources Regimbau (2011); Christensen (2019). Different types of sources can contribute to the SGWB, including compact binaries, core-collapse supernovae and isolated spinning neutron stars (e.g. Buonanno et al., 2005; Regimbau and Mandic, 2008; Crocker et al., 2015). Observing the SGWB will allow probing high-redshift populations, which are otherwise not accessible to ground-based detectors, although various methods are being developed to extract the SGWB from the detector noise in data from Advanced LIGO, Advanced Virgo and Kagra (LVK) Regimbau (2022); Renzini et al. (2022). Current upper limits for ground-based detectors are $\Omega_{\rm{GW}}(f)\leq 3.4\times 10^{-9}$ at $f=25$ Hz, where $\Omega_{\rm{GW}}$ is the dimensionless energy density of the SGWB Abbott, B. P. et al. (2017, 2018b); Abbott, B. P. et al. (2019, 2021); Abbott, R. et al. (2021). These upper limits contain valuable information on high-redshift populations of compact binaries, complementary to that of individual detections (e.g. Kowalska-Leszczynska et al., 2015; Dvorkin et al., 2016; Nakazato et al., 2016; Périgois et al., 2021; Cusin et al., 2019; Callister et al., 2020; Périgois et al., 2022; Lehoucq et al., 2023). For the LVK detectors, the dominant contribution to the SGWB is expected to be that of binary black holes. However, this will not be the case for ET and CE, since almost all the binary black hole mergers will be detected individually and subtracted from the data. The remaining SGWB will thus contain information on the unresolved BNS mergers.
In the case of the postmerger emission from a population of BNS mergers, we expect that if HMNS are formed, the characteristic GW frequencies will be enhanced in the SGWB due to the large number of unresolved sources contributing to the signal. This opens up the possibility to detect this EOS-dependent feature and to obtain complementary information on the postmerger stage. The goal of this Letter is to model the SGWB from a population of BNS, including the postmerger stage, and to highlight the existence of a so-far neglected additional and dominant component that can be detected with ET and CE.
BNS inspiral and postmerger waveforms— To construct a hybrid waveform, we combine in the frequency domain the analytic prescriptions referring to the inspiral and the postmerger. This combination takes place at a frequency $f_{\rm tr}$, where we move from the inspiral to the postmerger description (see End Matter for details). This frequency must be significantly smaller than the main frequencies appearing in the power spectral density (PSD) of the postmerger signal, i.e., $f_{1}$ and $f_{2}$, where $f_{1}<f_{2}$, so that no power from the postmerger phase is missed. Hereafter, for simplicity we set $f_{\rm tr}=f_{1}/2$, but the overall results depend only weakly on this assumption. The PSD in the $f_{1}$ frequency is significantly suppressed in the case on unequal-mass binaries but is here included for consistency with the use of a population of neutron-star binaries with equal-mass components. In practice, because the corresponding power is about one order of magnitude smaller than that in the $f_{2}$ peak, its impact in shaping the high-frequency component of the SGWB is very small.
The equation of state— In view of the uncertainties in the EOS and their effect on the waveform, we adopt an agnostic approach in this work. We use the findings of Refs. Altiparmak et al. (2022); Ecker and Rezzolla (2022, 2023); Jiang et al. (2023), who constructed more than $10^{7}$ EOSs consistent with constraints from fundamental physics and astronomical observations of neutron stars. Leveraging on this extensive dataset, it was possible to derive informative constraints on the relationship between the chirp mass $\mathcal{M}_{\mathrm{chirp}}$ and tidal deformability $\Lambda$ in terms of a probability density function (PDF) for these two quantities. A-priori, a single EOS does not correspond to a unique line in the $\mathcal{M}_{\mathrm{chirp}}$-$\Lambda$ plane because the same chirp mass may result from many combinations of BNS masses. However, when restricting the analysis to symmetric binaries, as we do throughout the rest of this work, this ambiguity disappears. In this case, for a given EOS, each chirp mass is uniquely associated to a single value of $\Lambda$. Therefore, extracting an EOS from a given distribution is equivalent to drawing a single value of $\Lambda$ from the PDF at each chirp mass. Similarly, extracting a certain value of $\Lambda$ from the allowed range is equivalent to singling-out a specific EOS.
Although the PDF found in Ref. Altiparmak et al. (2022) for a given chirp mass is not exactly a Gaussian, it is so to a first approximation [see Fig. (4) in Altiparmak et al. (2022)]. Hence, when assuming an underlining Gaussian distribution, we can relate the tidal deformability and the chirp mass (in solar masses) via a simple power-law of the type $\Lambda_{p}=\alpha+\beta\mathcal{M}^{\gamma}_{\mathrm{chirp}}$ where $\Lambda_{p}$ is normally distributed between the minimum and maximum values given by Altiparmak et al. (2022) $\Lambda_{\rm min\,(max)}=a+b\,\mathcal{M}_{\rm chirp}^{c}$ with $a=-50\,(-20)$, $b=500\,(1800)$, and $c=-4.5\,(-5.0)$. We here consider $\gamma=-3$, but we have checked that varying $\gamma$ leads to very little differences in SGWB, as the mass distribution of BNS has a clear main peak. In this way, we can build a set of fictitious $10^{3}$ EOSs such that the corresponding tidal deformability is normally distributed.
Once an EOS is set, we can compute the inspiral waveform. For the postmerger waveform, we need to know the characteristic frequencies $f_{1}$ and $f_{2}$ given the selected EOS and hence tidal deformability. Luckily, this is rather straightforward to do since universal relations have been found that provide analytic fits (see, e.g., Bauswein and Janka (2012); Takami et al. (2014a); Bauswein and Stergioulas (2015a); Rezzolla and Takami (2016)). Here, we employ the expressions for $f_{1}$ and $f_{2}$ proposed in Refs. Rezzolla and Takami (2016); Topolski et al. (2024) (and also employed in Bose et al. (2018)). We note that being only “quasi”-universal relations, these expressions come with some intrinsic uncertainties as estimated from the numerical simulations and the corresponding fits (see discussion in Topolski et al. (2024)). These uncertainties, however, do not play an important role here as we are interested in a SGWB that is naturally broad given the much larger uncertainties in the EOS.
Stochastic gravitational-wave background— The energy density in GWs per logarithmic unit of frequency and in units of the critical density of the Universe, $\rho_{\mathrm{c}}=3H_{0}/8\pi G$, is given by $\Omega_{\mathrm{GW}}:=({1}/{\rho_{c}})d\rho_{\mathrm{GW}}/{d}\ln{f}=({f}/{c\,\rho_{\mathrm{c}}})F(f)$, where $\rho_{\mathrm{GW}}$ is the GW energy density and $F(f)$ is the energy flux. When considering a population of $N$ sources, the energy flux per unit frequency is given by Allen and Romano (1999): $F(f):=({\pi c^{3}\,f^{2}}/{2G\,T})\sum_{i=1}^{N}[|\tilde{h}_{i}^{+}(f)|^{2}+|\tilde{h}_{i}^{\times}(f)|^{2}]$, where the index $i$ sums over the $N$ sources and $\tilde{h}_{i}^{+}(f)$ and $\tilde{h}_{i}^{\times}(f)$ are the Fourier transforms of the two GW polarization modes. The number of sources is taken to be proportional to the total observation time $T$. We have checked that our calculation converges at $T=1$ year and will use this value in the following (note that the background is stationary).
In order to construct our BNS catalogue, we follow the baseline_delays model from Ref. Lehoucq et al. (2023) and, more specifically, we assume all binaries are symmetric in mass and follow a Gaussian mass distribution with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. The BNS-merger rate is described by the star-formation rate from Ref. Vangioni et al. (2015) convolved with a distribution of delay times $t_{\mathrm{d}}$ that follows a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ Chruslinska et al. (2018). Using this model, we construct a catalogue containing about $N=3.8\times~{}10^{5}$ BNS that merge in the entire Universe during one year, consistent with other studies Iacovelli et al. (2022) (see Ref. Lehoucq et al. (2023)).
As we focus on the background, we must consider the flux contributions from non-detectable sources only. We assume that individual resolvable sources can be identified when their signal-to-noise ratio (SNR) exceeds a threshold $\mathrm{SNR}_{\mathrm{thr}}=12$  Meacher et al. (2015); Iacovelli et al. (2022), so that the contribution from these resolvable sources is removed by subtracting each individual signal, and we use the GWFish package Dupletsa et al. (2023) to compute the SNR. In this initial study we assume that the subtraction is noiseless, and leave more realistic approaches for future work Sachdev et al. (2020a); Regimbau et al. (2017); Sachdev et al. (2020b); Zhou et al. (2023).
Results— Figure 1 shows the SGWB from BNS mergers computed with our two descriptions for the modelling of the signal. More specifically, in the inspiral+prompt collapse model, hereafter “IPC” (blue-shaded areas), which is the one adopted so far in describing the SGWB from BNS mergers, the inspiral is assumed to be followed by a prompt collapse to a BH of the merger remnant. On the other hand, in the novel inspiral+postmerger model, hereafter “IPM”, proposed here (red-shaded areas), the signal includes the postmerger emission from a metastable HMNS that eventually collapses into a BH. For both models, we use ${\rm SNR}_{\mathrm{thr}}=12$, and the background was calculated using $10^{3}$ randomly sampled EOSs using the methodology described above. The shaded regions correspond to the $68$ and $95$ percentiles of the resulting distribution, respectively, while the solid lines indicate the median values in the distributions.
Figure 1: Energy density of the SGWB from BNS mergers, corresponding to either the IPC (blue-shaded areas) or the IPM (red-shaded areas) scenarios. Shown with different transparencies are the $68$ and $95$ percentile contours, while the solid lines report the medians. Also outlined are the power-integrated sensitivity curves of a single ET (dashed lines) and of a network composed of an ET detector and of two CE detectors (solid lines) relative to $1$ (grey lines) and $10$ (black lines) years of integration, respectively.
Figure 2: Top panel: The same as in Fig. 1 but where different colours show how $\Omega_{\rm GW}$ changes when assuming threshold SNR values of ${\rm SNR_{thr}}=12,20,40$, and $80$; note that $\tau_{2}=50\,{\rm ms}$ in all cases. Bottom panel: The same as in the top but when considering different values for the decay timescale $\tau_{2}=20,50$ and $250\,{\rm ms}$ (blue, green, and red, respectively). Note that ${\rm SNR_{thr}}=12$ in all cases.
As shown in Fig. 1, both models coincide at low frequencies (i.e., below $200\,{\rm Hz}$), which is expected since their inspiral parts are identical. They begin to diverge at approximately $350\,{\rm Hz}$. The power in the IPC model rapidly decreases as expected, since it does not include significant postmerger emission. In contrast, the IPM model exhibits a peak followed by a gradually declining plateau, and then experiences a rapid rise to a broad peak near $1.7\,{\rm kHz}$, which is about six times higher than the $350\,{\rm Hz}$ peak in $99\%$ of the EOSs considered. This peak is broad because of the redshift distribution of the sources and the intrinsic differences in the $10^{3}$ EOSs considered. We note that the high-frequency peak arises solely from the postmerger contribution, while the low-frequency peak is primarily influenced by the inspiral phase, as indicated by the overlap of the blue and red distributions. On the other hand, the width of the SGWB peak is due mostly to the redshift distribution of the sources, while the uncertainty in the EOSs contained in our sample is what leads to the variance shown with the shaded areas.
Black solid and dashed lines in Fig. 1 represent the power-law integrated sensitivity curves of two 3G detector configurations: a single ET configuration (dashed lines) and an admittedly optimistic configuration consisting of $2$ ET detectors, as well as two CE detectors (solid line), such type of network can drastically improve the background sensitivity Gupta, Ish et al. (2024); Reitze, David et al. (2019a); Maggiore et al. (2024). Sensitivity curves for both scenarios are shown either for one-year integration time (grey solid lines) or for ten years of observations (black lines) and with an SNR for SGWB detection of $\rho=1$.
The results in Fig. 1 clearly indicate that even after ten years, a single 3G detector will not be able to reach the peaks in the SGWB arising from the inspiral or from the postmerger. However, the network setup could detect the low-frequency peak with just one year of data and potentially the high-frequency peak with ten years of observation. Note that the improvement in sensitivity between the two scenarios is less pronounced than expected, primarily because ET and CE are geographically distant, leading to a suppressed overlap-reduction function at high frequencies, thereby limiting the potential gains.
We have so far assumed that all the individually detected sources are subtracted noiselessly from the data, leaving out the clear SGWB from unresolved BNS. In practice, however, this subtraction is challenging, especially for low-SNR sources, whose parameters are recovered with large uncertainties. To assess the impact that this subtraction has on the estimated SGWB, we consider scenarios in which the individual detection threshold is increased, which is equivalent to tagging as detected only the loudest sources. In this case, we expect better source subtraction, leading to less noise in the SGWB, which will also contain more sources.
We explore different detection thresholds in the left panel of Fig. 2, where we compare our default value $\mathrm{SNR}_{\mathrm{thr}}=12$ with higher values of $20,40$, and $80$. As expected, the amplitude of the SGWB increases with $\mathrm{SNR}_{\mathrm{thr}}$, which is encouraging as the background approaches the network sensitivity. Indeed, as we increase $\mathrm{SNR}_{\mathrm{thr}}$, fewer sources are identified as individual detections, and instead contribute to the background. At the same time, as $\mathrm{SNR}_{\mathrm{thr}}$ is increased, the positions of both the low- and high-frequency peaks shift to higher frequencies. This shift occurs because the background increasingly includes brighter sources, which are typically closer and less redshifted. Therefore, the peaks move to larger frequencies, where the detector sensitivity is lower, so that the actual detection of the SGWB at these frequencies will remain challenging despite the decrease in the background noise.
[…]
As expected, the PSD in the high-frequency peak of the postmerger signal increases linearly with the increase of $\tau_{2}$, thus considerably increasing the potential of detection of this part of the SGWB. At the same time, integration times longer than one year are needed to detect the high-frequency portion of the SGWB related to the postmerger. Interestingly, however, the increase in the decay time also leads to an increased power at lower frequencies, i.e., $500\lesssim f\lesssim 1000\,{\rm Hz}$, so that the detectability of the high-frequency portion of the SGWB is possible even if the maximum of the postmerger PSD is still below sensitivity. In turn, the results shown in the right panel of Fig. 2 also reveal that the detection of the high-frequency portion of the SGWB can provide precious information on the lifetime of the remnant and hence on the EOS. It is unlikely that stringent constraints on the EOS can be drawn from this measurement, which however will serve as an important consistency check. We will explore the implications of these results and how they can be used to infer the nuclear-matter EOS in a future work.
Conclusion— The SGWB generated by the inspiral and merger of binary neutron stars has so far been modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for very massive binaries, it is not what is expected to take place for systems having masses around the peak of the mass distribution of neutron-star binaries. Hence, the most likely outcome of the merger is that of producing a remnant that will survive for hundreds of milliseconds before collapse. During its short existence, the remnant can radiate enormous amounts of energy in GWs and, indeed, the radiated GW energy is 3-4 times larger than that emitted over the long inspiral.
These considerations have motivated us to include also the postmerger contribution to the PSD and hence calculate how the SGWB is modified by this important and dominant contribution. To this scope, we have implemented an analytic model of the postmerger GW signal that has been shown to provide a reasonably accurate description of the signal computed by numerical-relativity simulations Bose et al. (2018). The analytic waveform is dominated by the fundamental frequency of oscillation of the remnant $f_{2}/2$, that provides a narrow-banded emission around $f_{2}\simeq 1-2\,{\rm kHz}$. Quasi-universal relations have been employed to relate $f_{2}$ with the tidal deformability and the latter with the chirp mass of the binary. Finally, the analytic description of the GW emission and the parametric construction of the EOSs have then been combined with a model for the population of equal-mass BNS binaries Lehoucq et al. (2023). In this way, the postmerger contribution depends mostly on only two parameters: the decay timescale of the GW at the $f_{2}$ frequency, i.e., $\tau_{2}$, and the threshold value of the SNR at which a BNS can be resolved out of the background, i.e., ${\rm SNR_{thr}}$. Because both $\tau_{2}$ and ${\rm SNR_{thr}}$ are poorly known, in our analysis they are varied across the reasonable intervals in which they are expected to range.
Using this setup, we were able to compute the high-frequency postmerger contribution to the SGWB from about $\sim 3.8\times 10^{5}$ binaries (i.e., those that are expected to merge in the whole universe in one year Lehoucq et al. (2023)), finding that this indeed represents a new and dominant contribution. While being broad, this contribution peaks at $\sim 1-2\,{\rm kHz}$ and provides a contribution to the normalised GW energy density $\Omega_{\rm{GW}}\simeq 10^{-10}-10^{-9}$, hence much larger than what is expected when considering scenarios with the inspiral and collapse only. Clearly, being a background signal, the detectability of this contribution will also depend on the integration time employed and our analysis has shown that long integration times and a network of 3G detectors are likely to be needed for the detection. However, such a network setup could detect the peak of the high-frequency component in less than one year if $\tau_{2}\gtrsim 100\,{\rm ms}$. More importantly, the detection of the high-frequency portion of the SGWB can provide information that is either complementary or supplementary to that of the foreground signals. More specifically, the detection of the high-frequency component would constrain the typical lifetime of the remnant and since the latter depends on the EOS, it would provide complementary information on the EOS. On the other hand, when considering the reconstruction of the BNS population and its evolution with redshift, the detection of the high-frequency component – and specifically the location of the maximum and its broadness – will tell us about the whole population of binaries beyond what can be deduced from the loudest foreground sources. Overall, the results presented in this first analysis show that the postmerger contribution to the SGWB cannot be ignored as done so far and that it contains crucial information on the HMNS accessible with 3G detectors.
Notwithstanding the innovative nature of the procedure presented here, our approach can be improved in a number of ways. First, by considering that the decay time is not a simple mass dependant power-law, but actually a function of the EOS. Second, by extending the analysis to binaries that are not symmetric and hence introduce a degeneracy in the modelling of the chirp mass. Finally, by considering a more careful analysis of the foreground subtraction, which has been here treated simply in terms of the threshold SNR. We will explore these refinements of the model in future work.
Acknowledgements— This research is supported by the École Normale Supérieure Paris-Saclay CDSN PhD grant. This research is also supported in part by the ERC Advanced Grant “JETSET: Launching, propagation and emission of relativistic jets from binary mergers and across mass scales” (grant No. 884631), by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 “Strong-interaction matter under extreme conditions” – project number 315477589 – TRR 211, and by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). LR acknowledges the Walter Greiner Gesellschaft zur Förderung der physikalischen Grundlagenforschung e.V. through the Carl W. Fueck Laureatus Chair.
End Matter
On the energy radiated in GWs— A major focus of our work is to extend the contribution of the BNS merger signal to the SGWB by considering not only the inspiral part but also the postmerger emission. To appreciate simply why this is important and hence the relevance of our findings, we show in Fig. 3 the total GW energy radiated by a population of BNS when integrated over the entire frequency range ($E_{\mathrm{GW}}$) and normalized by the total energy considering only the inspiral phase ($E_{\mathrm{GW},0}$), as a function of the time before and after the merger $t-t_{\rm mer}$. Near the merger ($t-t_{\rm mer}\lesssim 0$) the energy radiated increases very rapidly, gaining an order of magnitude in the last $10\,{\rm ms}$ before the merger. Following the merger ($t-t_{\rm mer}>0$), the energy radiated continues to grow at a slower rate but more intensely, so that it can reach values much larger than those lost during the whole inspiral. Stated differently, BNS mergers are copious emitters of GWs during the inspiral, but it is the postmerger that provides the largest contribution to the radiated energy.
Of course, the exact amount of energy radiated will depend sensitively on the EOS describing the matter of which the BNS is composed of and which is essentially unknown. Instead of concentrating on a few specific EOSs, it is more convenient to consider a large ensemble of agnostic EOSs satisfying astrophysical, GW, or elementary-particle constraints Annala et al. (2018); Most et al. (2018); Greif et al. (2019); Altiparmak et al. (2022). This is what is represented with the blue-shaded band in Fig. 3, which reports the 99th percentile contour of an ensemble of $10^{3}$ EOSs generated using our EOS-agnostic approach. Similarly, the solid blue line reports the median of the distribution at each time in the evolution. In this way, it is straightforward to appreciate that since most of the energy is lost after the merger, it is important to include also the postmerger emission when calculating the BNS GW background (similar considerations apply when considering the radiated angular momentum Ecker et al. (2025); Bernuzzi et al. (2016)).
Figure 3: Evolution of the radiated GW energy when normalized to the at the merger. Shown with blue- and red-shaded areas are the corresponding values during the inspiral (blue) and the postmerger. The variance in the shaded regions marks the $99\%$ percentile.
Modelling the BNS waveforms— To model the inspiral part of the GW waveform, we use the IMRPhenomD_NRTidalv2 model Dietrich, Tim et al. (2019); Dietrich et al. (2019) from the LALSuite simulation package LIGO Scientific Collaboration (2020). This model extends the IMRPhenomD waveform Husa et al. (2016); Khan et al. (2016), designed for binary black hole systems, by incorporating EOS-dependent tidal effects. Specifically, the template uses the dimensionless tidal deformability parameter $\Lambda:=G^{4}\lambda/c^{5}M^{5}$, where $M$ is the stellar mass and the tidal deformability parameter $\lambda$ is defined as the ratio between the induced quadrupole and the external gravitational field ($G$ and $c$ are the gravitational constant and the speed of light, respectively). The model assumes that the merger remnant promptly collapses to a black hole, thus leading to very little power in the postmerger phase Dietrich et al. (2019). As commented above, this is a very drastic assumption and applicable only to a rather restricted set of binaries with total gravitational mass that is roughly given by $M_{\rm tot}\gtrsim 1.45\,M_{{}_{\rm TOV}}\,\sim 3.38\,M_{\odot}$ Koeppel et al. (2019), where $M_{{}_{\rm TOV}}$ is the maximum nonrotating gravitational mass for a given EOS and the estimate comes from assuming a maximum mass $M_{{}_{\rm TOV}}\sim 2.33\,M_{\odot}$ Margalit and Metzger (2017); Rezzolla et al. (2018) (see also Bauswein et al. (2013); Tootle et al. (2021); Kölsch et al. (2022); Kashyap et al. (2022) for a discussion on the threshold mass in BNS mergers).
Numerical-relativity simulations suggest that BNS mergers typically produce a metastable HMNS, which persists for several tens of milliseconds before ultimately collapsing into a spinning black hole Baiotti and Rezzolla (2017); Paschalidis (2017). Recent studies showed that the GW signal emitted by this HMNS exhibits several prominent peaks in the frequency domain Stergioulas et al. (2011); Bauswein and Janka (2012); Bauswein and Stergioulas (2015a); Rezzolla and Takami (2016); Takami et al. (2014a, 2015a); Topolski et al. (2024). Here, we use the results of Ref. Bose et al. (2018), which presents analytical fits for the postmerger GW signal based on an extensive catalogue of time-domain numerical waveforms. These fits express the time-domain strain as a sum of damped sinusoids with a time-evolving instantaneous frequency as $h_{+}(t)=h_{0}\,\mathrm{e}^{-t/\tau_{1}}[\mathrm{sin}(2\pi f_{1}t)+\mathrm{sin}(2\pi(f_{1}-f_{1\epsilon})t)+\mathrm{sin}(2\pi(f_{1}+f_{1\epsilon})t)]+\mathrm{e}^{-t/\tau_{2}}\,\mathrm{sin}(2\pi f_{2}t+2\pi\gamma_{2}t^{2}+2\pi\xi_{2}t^{3}+\pi\beta_{2})$, where $h_{0}=G\mathcal{M}_{\mathrm{chirp}}/(2Dc^{2})$ with $D$ the distance to the source, $t=0$ is the merger time, the frequencies $f_{1}$ and $f_{2}$ represent the dominant postmerger GW modes with $f_{2}>f_{1}$, $\tau_{1}$ and $\tau_{2}$ are the decay times of each mode, and $f_{1\epsilon}$, $\beta_{2}$, $\gamma_{2}$ and $\xi_{2}$ are constants Stergioulas et al. (2011); Takami et al. (2014b); Rezzolla and Takami (2016). The first peak, $f_{1}$, is short-lived (with a decay time $\tau_{1}$ of the order of a few ms) and most probably originates from damped collisions of the two stellar cores immediately after the merger Takami et al. (2015b). The second peak, $f_{2}$, corresponds to the spin frequency of the deformed HMNS and has a decay time $\tau_{2}$ of the order of several tens of milliseconds Takami et al. (2014b); Bauswein and Stergioulas (2015b). However, the expected values, and mass dependencies of both $\tau_{1}$ and $\tau_{2}$ are still rather uncertain as they sensitively depend on the EOS and the microphysical effects taking place after the merger. Interestingly, when considering the information collected with the GW170817 event, some studies have constrained the maximum lifetime of the remnant from that event to approximately one second Gill et al. (2019); Murguia-Berthier et al. (2021). Overall, it is clear that the lifetime of the HMNS can vary significantly between a few to several thousands of milliseconds.
[…]
Figure 4 illustrates the construction of the hybrid waveform for a BNS system with masses $M_{1}=M_{2}=1.25\,M_{\odot}$, at a distance from observer $D=50\,{\rm Mpc}$ when employing GNH3 EOS discussed in Ref. Bose et al. (2018). Shown with a blue dashed line is the inspiral-and-merger signal followed by a prompt collapse and described by the IMRPhenomD_NRTidalv2 model, the postmerger signal with a black dashed line, and the combined (hybrid) signal with a red solid line. We also show with a solid green line the ET strain sensitivity. The transition frequency $f_{\rm tr}$ is meant to mark therefore the boundary between the grey and white regions. In the grey region, the hybrid waveform is simply given by the IMRPhenomD_NRTidalv2 waveform. Note that in the white region, the hybrid waveform is given by the sum of the inspiral and postmerger strains and that the latter is negligible below $f_{\rm tr}$, thus ensuring no significant power is lost or added. To properly combine the two waveforms above $f_{\rm tr}$, we must ensure phase coherence, which is not guaranteed a priori, as we have not explicitly computed the temporal phase $\beta_{2}$. Therefore, at $f_{\rm tr}$, we adjust the phase of the IMRPhenomD_NRTidalv2 waveform to match the postmerger one, thus ensuring phase continuity.
Figure 4: Evolution in the frequency-domain of the GW strain for a BNS with masses $m_{1}=m_{2}=1.25\,M_{\odot}$ at a distance of observer $D=50\,{\rm Mpc}$ and relative to the GNH3 EOS Bose et al. (2018). The IMRPhenomD_NRTidalv2 waveform is shown with a black dashed line, the HMNS postmerger with a blue dashed line, and the combined waveform with a red solid line. The frequency $f_{\rm tr}$ marks the transition between the inspiral (grey-shaded are) and the postmerger, while the green solid curve reports the sensitivity of ET. Also reported are the main components $f_{2}$ and $f_{1}$ of the spectrum of equal-mass binaries.
Modelling the BNS population— In order to calculate the SGWB from the entire population of unresolved BNS mergers, we need to model the merger rate as a function of redshift and system parameters. Here, we consider (equal-mass) symmetric non-eccentric binaries, so that each BNS is characterized by the component masses $m_{1}=m_{2}=m$, tidal deformability parameters $\Lambda_{1}=\Lambda_{2}=\Lambda$, redshift $z$, sky position (right ascension $\alpha$ and declination $\delta$), orbital inclination angle $\iota$, polarization angle ($\psi$), coalescence phase $\Phi_{\mathrm{c}}$ and coalescence time set to be $t_{\mathrm{c}}=0$. The mass distribution we used is derived from Galactic observations and is assumed to be independent of redshift. According to Ref. Farrow et al. (2019), this distribution is well-modeled by a Gaussian with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. Both neutron stars in the binary are assumed to follow this distribution, which is tightly concentrated around the mean, making it a non-critical parameter for this study.
We note that this mass distribution may depend on the formation scenario of the binary. More specifically, the mass distribution of BNS detected via GW may differ from that of Galactic binaries, potentially explaining the higher total mass observed in the BNS merger GW190425 collaboration (2020); Romero-Shaw et al. (2020). Furthermore, the recent GW190814 event suggests a black hole-neutron star (BHNS) merger involving a neutron star significantly heavier than typical Galactic counterparts Abbott, R. et al. (2020), although there are arguments favouring a black-hole origin for the secondary Nathanail et al. (2021).
We assume that the BNS merger rate follows the star formation rate (SFR) convolved with a distribution of delay times. Specifically, we adopt the baseline_delays model from Ref. Lehoucq et al. (2023), where the merger rate is defined as: $R_{\mathrm{merg}}(t)=R_{0}\int^{t_{\mathrm{d},\mathrm{max}}}_{t_{\mathrm{d},\mathrm{min}}}\mathrm{SFR}(t-t_{\mathrm{d}})\,P(t_{\mathrm{d}})\,\mathrm{d}t_{\mathrm{d}}$. Here the SFR is taken from Ref. Vangioni et al. (2015) and $P(t_{\mathrm{d}})$ is the distribution of time intervals $t_{\mathrm{d}}$ between the formation of stellar progenitors and the eventual BNS merger. We here assume a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ Chruslinska et al. (2018), and the delay times to range from $t_{\mathrm{d},\mathrm{min}}=10\,{\rm Myr}$ to $t_{\mathrm{d},\mathrm{max}}$ set by the Hubble time. The parameter $R_{0}$ is a normalization constant that calibrates the merger rate at $z=0$ to the local BNS merger rate inferred from the GWTC-3 LIGO-Virgo catalogue: $R_{0}=44^{+96}_{-34}~{}~{}{\rm yr}^{-1}\,{\rm Gpc}^{-3}$ Abbott, R. et al. (2023).